e the following exchange quotes to help answer problems 18-19. NYSE NASDAQ Bid Ask Bid Ask CompanyX CompanyY $ 149.55 150.45 149.87 150.13 y YS 127.88 S 128.12 127.70 S 128.30 18. Currently you have $120.00 in cash and 1000 shares of Stock Y in your brokerage account. You plan to sell your 1000 shares of Stock Y and use the proceeds to purchase as many shares of Stock X as possible. Assume you cannot borrow or buy partial shares of stock: you will also buy or sell at the best price offered. Based on the quotes provided, the cash balance in your brokerage account after you purchase the shares of Stock X is closest to: A. $(60.89) B. $59.11 C. $87.95 D. $89.24 E. $106.10 19. Based on the bid-ask spreads for Stock x & Y, the NASDAQ is more liquid than the NYSE while Stock Y is more liquid than Stock X. Evaluate the underlined words in italics. True or False? Hint: Calculate spread as a percentage of the midpoint of the bid-ask quotes A. True B. False 20. During lecture we compared Warren Buffets Berkshire Hathaway (BRK.A), which has a market cap of around $480B, to Amazon (AMZN), which has a market cap of around $790B. We concludes True or False? A. True of a stock, the greater the stocks liquidity, Evaluate the underlined words in italics. er B. False g lecture we compared the stock market to the bond market. Which of the following statements (- Ill) is (are) most likely FALSE: I. Stocks are generally more liquid than bonds ll. There is more money invested in the US stock market than the US bond market ll. The fact that bonds are priced relatively infrequently made things worse during the housing crisis A. l D. I, ll, lll E. None of the statements
Answer #1

18) HE HAVE 1000 SHARE OF COMPANY Y = 127.7*1000 = (127700)

SELL 1000 SHARE OF COMPANY Y @128.30 = 128.30*1000 = (128300)

PROFIT = 127700-128300 = (600)

CURRENT CASH IN HAND = 127700+600+120= 128420

WE CAN BUY = 858.71 SHARES OF STOCK X (128420/149.55)

BUT WE CANT BUY SHARE IN DECIMAL SO THE CASH WE HAVE IN HAND IS 0.71*149.55 = (106.18)

OPTION “E” IS THE CORRECT ANSWER FOR QUESTION 18

QUESTION 19)

NYSE

NASDAQ

BID

ASK

BID

ASK

COMPANY X

149.55

150.45

149.87

150.13

COMPANY Y

127.88

128.12

127.7

128.3

SPREAD

NYSE

NASDAQ

COMPANY X

0.602%

0.173%

COMPANY Y

0.188%

0.470%

FORMULA

(ASK-BID)/BID*100

ANSWER 19) FALSE

QUESTION 20)

Generally Large Cap Company will have more liquidity

Answer 20) TRUE

ANS 21) E